# bibCyclic.bib

### v2

@@ 6.3 Matrix elements of the fusion operator
We recall the following definition from [OkSm].
@@
Let $A\in\text{End}(\mathsf{F}\otimes\mathsf{F})$ be an operator of the form $A=\bigoplus_{\alpha\in\mathbb{Z}^{l}}A_{\alpha}$ where for any $\mathsf{v}$ and $\mathsf{v}^{\prime}$, $A_{\alpha}(K_{\mathsf{T}}(\mathcal{M}(\mathsf{v},\delta_{0})\times\mathcal{M}(\mathsf{v}^{\prime},\delta_{0})))\subset K_{\mathsf{T}}(\mathcal{M}(\mathsf{v}+\alpha,\delta_{0})\times\mathcal{M}(\mathsf{v}^{\prime}-\alpha,\delta_{0}))$. Then $A$ is upper triangular if $A_{\alpha}=0$ unless $\sum_{i=0}^{l-1}\alpha_{i}\geq 0$. Furthermore, if $A_{0}=1$, then $A$ is strictly upper triangular.
Let $F=\mathcal{M}(\mathsf{v},\delta_{0})\times\mathcal{M}(\mathsf{v}^{\prime},\delta_{0})$ so that $K_{\mathsf{T}}(F)$ is a direct summand of $\mathsf{F}\otimes\mathsf{F}$. Let $\Omega$ be the $K_{\mathsf{T}}(pt)$ linear operator on $\mathsf{F}\otimes\mathsf{F}$ which acts on $K_{\mathsf{T}}(F)$ by multiplication by one fourth the codimension of $F$ in $\mathcal{M}(\mathsf{v}+\mathsf{v}^{\prime},2\delta_{0})$. An explicit formula for $\Omega$ is given in section 2.3.7 of [OkSm]. Let $Z_{(1)}$ be the operator which acts on $K_{\mathsf{T}}(F)$ as multiplication by $\prod_{i=0}^{l-1}z_{i}^{\mathsf{v}_{i}}$.
Proposition 6.5 ([OkSm] Proposition 7).
There is a unique element $J(z)\in(\rho_{f}\otimes\rho_{f})(h(\mathscr{U}_{\hbar}(\dot{\mathfrak{g}}))\otimes h(\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})))$ such that $J(z)$ is strictly upper triangular and solves the ABRR equation:
$(\rho_{f}\otimes\rho_{f})(h(\mathcal{R}))Z^{-1}_{(1)}J(z)Z_{(1)}=\hbar^{-\Omega}J(z)\hbar^{\Omega}$
@@
The operator $J(z)$ is known as the “fusion operator”. It is defined in [OkSm] using the wall 00 $R$-matrix. Due to Theorem 6.1, we can equivalently use the $R$-matrix of the horizontal subalgebra.
Our interest in $J(z)$ is due to Proposition 7.6 below. The general formula for the solution to these equations is complicated, see [zhu1]. For our purposes, we will only need certain matrix elements of $J(z)$ for which the formulas are much simpler. We first consider matrix elements of $h(\mathcal{R})$ in the basis of wreath Macdonald polynomials.
@@
$\langle H_{\nu}\otimes 1|(\rho_{f}\otimes\rho_{f})h(\mathcal{R})|H_{\lambda}\otimes H_{\mu}\rangle=\langle H_{\nu}\otimes 1|(\rho_{f}\otimes\rho_{f})h(\mathcal{R}_{0})|H_{\lambda}\otimes H_{\mu}\rangle$
@@
Recall that, under the identification of Theorem 3.4, $(\rho_{f}\otimes\rho_{f})(h(\mathcal{R}))=(\rho_{b}\otimes\rho_{b})(\omega\otimes\omega)h(\mathcal{R})=(\rho_{b}\otimes\rho_{b})v(\mathcal{R})$, where $\omega$ is the Miki automorphism. So it suffices to consider the $R$-matrix of the vertical subalgebra in the bosonic representation, the formulas for which are given in Theorem 3.9 of [Wen]. In particular, the coefficients of the $e_{i}$ (resp. $f_{i}$) current in the vertical subalgebra act on the bosonic representation in the basis of wreath Macdonald polynomials by removing (resp. adding) a box of content $i$ from the core.
Now we apply the KT-factorization: $\mathcal{R}=\mathcal{R}^{+}\mathcal{R}_{0}\mathcal{R}^{-}$. Recall from the discussion at the beginning of section 6.2 that the second tensorands of $\mathcal{R}^{+}$ involve only coefficients of the $f$ currents. Since it is not possible to obtain the vacuum by adding a box to a partition, we deduce that $\langle H_{\nu}\otimes 1|\mathcal{R}^{+}|v\rangle=\langle H_{\nu}\otimes 1|v\rangle$ for any $v\in\mathsf{F}\otimes\mathsf{F}$.
Recall also from the beginning of section 6.2 that the second tensorands of $\mathcal{R}^{-}$ involve only the coefficients of the $e_{i}$ currents (for $i\neq 0$). Since these coefficients can never remove a box of content 0 from the core, we deduce that $\mathcal{R}^{-}(H_{\lambda}\otimes H_{\mu})=H_{\lambda}\otimes H_{\mu}+\ldots$, where the dots stand for a linear combination of wreath Macdonald polynomials $H_{\gamma}\otimes H_{\delta}$ where $\text{core}_{l}(\delta)\neq\emptyset$.
Finally, by Lemma 6.3, $\mathcal{R}^{0}$ preserves the $l$-core of each tensorand. The result follows.
@@
$\langle H_{\nu}\otimes 1|(\rho_{f}\otimes\rho_{f})J(z)|H_{\lambda}\otimes H_{\mu}\rangle=\langle H_{\nu}\otimes 1|(\rho_{f}\otimes\rho_{f})J_{0}(z)|H_{\lambda}\otimes H_{\mu}\rangle$
@@
By Proposition 6.7, the matrix elements in the statement of the proposition agree with those of the unique strictly upper triangular solution $J_{0}(z)$ of the equation
$(\rho_{f}\otimes\rho_{f})(h(\mathcal{R}_{0}))Z^{-1}_{(1)}J_{0}(z)Z_{(1)}=\hbar^{-\Omega}J_{0}(z)\hbar^{\Omega}$ (10)
restricted to the subspace $\mathsf{F}_{\emptyset}\otimes\mathsf{F}_{\emptyset}\subset\mathsf{F}\otimes\mathsf{F}$ where $\mathsf{F}_{\emptyset}$ is the subspace spanned by wreath Macdonald polynomials with empty $l$-core.
From [OkSm] formula (35), $\Omega|_{K_{\mathsf{T}}(\mathcal{M}(n_{1})\times\mathcal{M}(n_{2}))}=\frac{1}{2}(n_{1}+n_{2})$.
To solve the previous equation, we make the ansatz inspired by Lemma 6.3:
@@
$J^{(i)}_{n}(z)z^{-n}+\hbar^{n/2}-\hbar^{-n/2}=J^{(i)}_{n}(z)$
where we have written $z$ for $\prod_{i=0}^{l-1}z_{i}$. Solving gives
$J^{(i)}_{n}(z)=\frac{(\hbar^{n/2}-\hbar^{-n/2})}{1-z^{-n}}$
@@
Note that for $l=1$ we recover the fusion operator for $\text{Hilb}^{n}(\mathbb{C}^{2})$, see section 7 of [OkSm].
@@ 7.1 Limit of vertex function
Let $\mathcal{M}=\mathcal{M}(\mathsf{v},\mathsf{w})$ be a $\hat{A}_{l-1}$ quiver variety. Split the framing as $\mathsf{w}=u_{1}\mathsf{w}^{\prime}+u_{2}\mathsf{w}^{\prime\prime}$ and let the corresponding subtorus of the framing torus be $\mathsf{A}^{\prime}$. We consider the limit of the vertex $V^{\left(\tau\right)}_{\mathcal{M}}$ of $\mathcal{M}$ as $u_{2}\to\infty$. It is proven in Lemma 7.3.5 of [Ok17] that $\lim_{u_{2}\to\infty}V^{\left(\tau\right)}_{\mathcal{M}}$ factorizes, up to a shift, as long as the limit of the descendant $\tau$ exists. Here we make the shift explicit.
@@
We denote $F^{\text{st}}=\mathcal{M}^{\text{st}}(\mathsf{v}^{\prime},\mathsf{w}^{\prime})\times\mathcal{M}^{\text{st}}(\mathsf{v}^{\prime\prime},\mathsf{w}^{\prime\prime})$. Hence if $\tau\in\mathcal{D}(\mathsf{v},\mathsf{w})$, then $\tau|_{F^{\text{stack}}}\in\mathcal{D}(\mathsf{v}^{\prime},\mathsf{w}^{\prime})\otimes\mathcal{D}(\mathsf{v}^{\prime\prime},\mathsf{w}^{\prime\prime})\otimes\mathbb{Q}[u_{1}^{\pm 1},u_{2}^{\pm 1}]$.
@@
This is Lemma 7.3.5 of [Ok17] spelled out for cyclic type $A$ quiver varieties. In general, the powers of $q$ arises from the polarization, and the powers of $-\hbar^{1/2}$ are the components of the vector $\mathsf{w}^{\prime\prime}+C\mathsf{v}^{\prime\prime}$ and $-\mathsf{w}^{\prime}-C\mathsf{v}^{\prime}$, where $C$ is the Cartan matrix of the quiver. ∎
@@
If, in particular, $\mathcal{M}=\mathcal{M}((n,n,\ldots,n),2\delta_{0})$, $\mathcal{M}^{\prime}=\mathcal{M}(n)$, and ′′$\mathcal{M}^{\prime\prime}=\mathcal{M}(0)=\{pt\}$, then
@@ 7.2 Limit of capping operator
Consider $\mathcal{M}:=\mathcal{M}(\mathsf{w})$ and split the framing as $\mathsf{w}=u_{1}\mathsf{w}^{\prime}+u_{2}\mathsf{w}^{\prime\prime}$. Let the corresponding subtorus of the framing torus be $\mathsf{A}^{\prime}$. Then $\mathcal{M}(\mathsf{w})^{\mathsf{A}^{\prime}}=\mathcal{M}(\mathsf{w}^{\prime})\times\mathcal{M}(\mathsf{w}^{\prime\prime})=:\mathcal{M}^{\prime}\times\mathcal{M}^{\prime\prime}$. Let $\text{Stab}:K_{\mathsf{T}}(\mathcal{M}(\mathsf{w})^{\mathsf{A}^{\prime}})\to K_{\mathsf{T}}(\mathcal{M}(\mathsf{w}))$ be the stable envelope [Ok17] with the chamber such that $a=u_{1}/u_{2}$ is an attracting weight, polarization given by
@@
Smirnov’s paper [Sm16] considers only the Jordan quiver. Nevertheless, his techniques apply equally well to all quivers, with the only needed modification being the fixed-component dependent shift in the statement of the theorem.
@@
This follows immediately from the three axioms for stable envelopes, see [OkSm]. ∎
@@
Let $\mathscr{L}_{0}^{(2)}$ be the operator of tensor multiplication by the 00th tautological line bundle of $\mathcal{M}(2\delta_{0})$. The component of
@@ 8 Derivation of main formula
We now have all the ingredients needed to prove the main theorem:
@@
Let
and
Then
@@
We first set up the notations for the proof. Let $\mathcal{M}^{(2)}_{\emptyset}=\bigsqcup_{n}\mathcal{M}((n,n,\ldots,n),2\delta_{0})$ and let $\mathsf{A}$ be the (rank two) framing torus, which has coordinates $u_{1}$ and $u_{2}$. The superscript $(2)$ refers to the size of the framing, while the subscript $\emptyset$ refers to the fact that the $K$-theory of $\mathcal{M}^{(2)}_{\emptyset}$ has a basis given by wreath Macdonald polynomials with empty l-core. Let $\tau^{(2)}_{0}$ denote the descendant from Definition 4.6 on $\mathcal{M}^{(2)}_{\emptyset}$ corresponding to the 00th tautological bundle. Let $V^{\left(\tau^{(2)}_{0}\right)}$ denote the corresponding descendant vertex function. Let $\Psi^{(2)}$ denote the capping operator for $\mathcal{M}^{(2)}_{\emptyset}$. Let $\mathscr{L}_{0}^{(2)}$ denote the operator of tensor multiplication in $K$-theory by the 00th tautological line bundle of $\mathcal{M}^{(2)}_{\emptyset}$.
@@
${K_{\mathsf{T}}\left(\mathcal{M}^{(2)}_{\emptyset}\right)}$${K_{\mathsf{T}}\left((\mathcal{M}^{(2)}_{\emptyset})^{\mathsf{A}}\right)}$${K_{\mathsf{T}}\left(\mathcal{M}^{(1)}_{\emptyset}\times\mathcal{M}^{(1)}_{\emptyset}\right)}$$\scriptstyle{\text{Res}_{\emptyset}}$
@@
$\sum_{\begin{subarray}{c}\lambda\\ \text{core}_{l}(\lambda)=\emptyset\end{subarray}}N_{\lambda}^{-1}H_{\lambda}[x]\hat{V}^{\left(\tau_{0}\right)}_{\lambda}|_{z_{0}=z_{0}(-\hbar^{1/2})}\\ =\left(J(z)^{-1}\sum_{\begin{subarray}{c}\lambda,\mu\\ \text{core}_{l}(\lambda)=\emptyset\\ \text{core}_{l}(\mu)=\emptyset\end{subarray}}N_{\lambda}^{-1}N_{\mu}^{-1}H_{\lambda}[x]H_{\mu}[y]\tau_{0}|_{\lambda}\right)\bigg{|}_{y=0}$ (15)
From Theorem 6.8, $J(z)^{-1}|_{K_{\mathsf{T}}(\mathcal{M}^{(1)}_{\emptyset})\otimes K_{\mathsf{T}}(\mathcal{M}^{(1)}_{\emptyset})}$, followed by $y=0$, acts by $p^{(i)}_{n}[x]\mapsto p^{(i)}_{n}[x]$ and $p^{(i)}_{n}[y]\mapsto p^{(i)}_{n}[y]+\frac{(\hbar^{-n/2}-\hbar^{n/2})}{1-z^{-n}}p^{(i)}_{n}[x]$. Combining this with Lemma 2.8 shows that the right side of (15) is equal to
$\left(J(z)^{-1}\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left((1-(-u)^{n})p^{(i)}_{n}[x]+(-u\hbar^{1/2})^{n}p^{(i)}_{n}[y]\right)\right)\right)\Bigg{|}_{y=0}\\ =\left(\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left((1-(-u)^{n})p^{(i)}_{n}[x]+p^{(i)}_{n}[y]+\frac{(-u)^{n}(1-\hbar^{n})}{1-z^{-n}}p^{(i)}_{n}[x]\right)\right)\right)\Bigg{|}_{y=0}\\ =\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}p^{(i)}_{n}[x]}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left((1-(-u)^{n})+\frac{(-u)^{n}(1-\hbar^{n})}{1-z^{-n}}\right)\right)$
@@
$\sum_{\begin{subarray}{c}\lambda\\ \text{core}_{l}(\lambda)=\emptyset\end{subarray}}N_{\lambda}^{-1}H_{\lambda}[x]\hat{V}^{\left(\tau_{0}\right)}_{\lambda}=\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}p^{(i)}_{n}[x]}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left(1-\frac{(-\hbar^{1/2}u)^{n}-(\hbar uz)^{n}}{(\hbar^{1/2})^{n}-(-z)^{n}}\right)\right)$
@@
[…]
Jeffrey Ayers
University of North Carolina
Chapel Hill, NC
jeff97@live.unc.edu
Hunter Dinkins
Northeastern University
Boston, MA
h.dinkins@northeastern.edu

### v3

@@ 6.3 Matrix elements of the fusion operator
We recall the following definition from [48].
@@
Let $A\in\text{End}(\mathsf{G}\otimes\mathsf{G})$ be an operator of the form $A=\bigoplus_{\alpha\in\mathbb{Z}^{l}}A_{\alpha}$ where for any $\mathsf{v}$ and $\mathsf{v}^{\prime}$, $A_{\alpha}(K_{\mathsf{T}}(\mathcal{M}(\mathsf{v},\delta_{0})\times\mathcal{M}(\mathsf{v}^{\prime},\delta_{0})))\subset K_{\mathsf{T}}(\mathcal{M}(\mathsf{v}+\alpha,\delta_{0})\times\mathcal{M}(\mathsf{v}^{\prime}-\alpha,\delta_{0}))$. Then $A$ is upper triangular if $A_{\alpha}=0$ unless $\sum_{i=0}^{l-1}\alpha_{i}\geq 0$. Furthermore, if $A_{0}=1$, then $A$ is strictly upper triangular.
Let $F=\mathcal{M}(\mathsf{v},\delta_{0})\times\mathcal{M}(\mathsf{v}^{\prime},\delta_{0})$ so that $K_{\mathsf{T}}(F)$ is a direct summand of $\mathsf{G}\otimes\mathsf{G}$. Recall that $\mathsf{G}\cong\mathsf{B}$ has a basis given by wreath Macdonald polynomials. For later use, we let $\mathsf{G}_{\emptyset}$ be the subspace spanned by wreath Macdonald polynomials with empty $l$-core.
Let $\Omega$ be the $K_{\mathsf{T}}(pt)$ linear operator on $\mathsf{G}\otimes\mathsf{G}$ which acts on $K_{\mathsf{T}}(F)$ by multiplication by one fourth the codimension of $F$ in $\mathcal{M}(\mathsf{v}+\mathsf{v}^{\prime},2\delta_{0})$. An explicit formula for $\Omega$ is given in section 2.3.7 of [48]. Let $Z_{(1)}$ be the operator which acts on $K_{\mathsf{T}}(F)$ as multiplication by $\prod_{i=0}^{l-1}z_{i}^{\mathsf{v}_{i}}$.
Proposition 6.7 ([48] Proposition 7).
There is a unique element $J(z)\in(\rho_{g}\otimes\rho_{g})(h(\mathscr{U}_{\hbar}(\dot{\mathfrak{g}}))\otimes h(\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})))$ such that $J(z)$ is strictly upper triangular and solves the ABRR equation:
$(\rho_{g}\otimes\rho_{g})(h(\mathcal{R}))Z^{-1}_{(1)}J(z)Z_{(1)}=\hbar^{-\Omega}J(z)\hbar^{\Omega}$
@@
The operator $J(z)$ is known as the “fusion operator”. It is defined in [48] using the wall 0 $R$-matrix. Due to Theorem 6.1, we can equivalently use the $R$-matrix of the horizontal subalgebra.
Our interest in $J(z)$ is due to Proposition 7.6 below. The general formula for the solution to these equations is complicated, see [67]. For our purposes, we will only need certain matrix elements of $J(z)$ for which the formulas are much simpler. We first consider matrix elements of $h(\mathcal{R})$ on $\mathsf{G}$ in the basis of torus fixed points.
@@
$\langle[I_{\nu}]\otimes 1|(\rho_{g}\otimes\rho_{g})h(\mathcal{R})|[I_{\lambda}]\otimes[I_{\mu}]\rangle=\langle[I_{\nu}]\otimes 1|(\rho_{g}\otimes\rho_{g})h(\mathcal{R}_{0})|[I_{\lambda}]\otimes[I_{\mu}]\rangle$
@@
We apply the KT-factorization: $\mathcal{R}=\mathcal{R}^{+}\mathcal{R}_{0}\mathcal{R}^{-}$. Recall the discussion at the beginning of section 6.2 regarding our choice of normal order of the affine roots. By (12) and Lemma 6.4, in the fixed point basis of $\mathsf{G}$, $(\rho_{g}\otimes\rho_{g})h(\mathcal{R}^{+})=1+\ldots$ where the dots represent terms that remove boxes of content other 0 from the second tensorand. Since any nonempty partition has a box of content 0, we deduce that $\langle[I_{\nu}]\otimes 1|(\rho_{g}\otimes\rho_{g})h(\mathcal{R}^{+})|v\rangle=\langle[I_{\nu}]\otimes 1|v\rangle$ for any $v\in\mathsf{G}\otimes\mathsf{G}$.
Again, by (13) and Lemma 6.4, we see that $(\rho_{g}\otimes\rho_{g})h(\mathcal{R}^{-})=1+\ldots$ where the dots stand for terms that remove boxes from the second tensorand but cannot remove an equal number of boxes of each color. Thus if $\lambda$ and $\mu$ have empty $l$-core, then $(\rho_{g}\otimes\rho_{g})h(\mathcal{R}^{-})([I_{\lambda}]\otimes[I_{\mu}])=[I_{\lambda}]\otimes[I_{\mu}]+\ldots$, where the dots stand for a linear combination of classes $I_{\gamma}]\otimes[I_{\delta}]$ where $\text{core}_{l}(\delta)\neq\emptyset$.
Finally, by Lemma 6.5, $(\rho_{g}\otimes\rho_{g})h(\mathcal{R}^{0})$ preserves the $l$-core of each tensorand. The result follows.
Recall the identification $\mathsf{G}\cong\mathsf{B}$ which maps $I_{\lambda}]$ to $H_{\lambda}$.
@@
$\langle[I_{\nu}]\otimes 1|J(z)|[I_{\lambda}]\otimes[I_{\mu}]\rangle=\langle H_{\nu}\otimes 1|J_{0}(z)|H_{\lambda}\otimes H_{\mu}\rangle$
@@
By Proposition 6.9, the matrix elements in the left side of the statement of the proposition agree with those of the unique strictly upper triangular solution $J_{0}(z)$ of the equation
$(\rho_{g}\otimes\rho_{g})(h(\mathcal{R}_{0}))Z^{-1}_{(1)}J_{0}(z)Z_{(1)}=\hbar^{-\Omega}J_{0}(z)\hbar^{\Omega}$ (14)
restricted to the subspace $\mathsf{G}_{\emptyset}\otimes\mathsf{G}_{\emptyset}\subset\mathsf{G}\otimes\mathsf{G}$.
From [48] formula (35), $\Omega|_{K_{\mathsf{T}}(\mathcal{M}(n_{1})\times\mathcal{M}(n_{2}))}=\frac{1}{2}(n_{1}+n_{2})$.
To solve the previous equation, we use the identification $\mathsf{G}\cong\mathsf{B}$ and make the ansatz inspired by Lemma 6.5:
@@
$J^{(i)}_{n}(z)(z_{0}\ldots z_{l-1})^{-n}+\hbar^{n/2}-\hbar^{-n/2}=J^{(i)}_{n}(z)$
Solving gives
$J^{(i)}_{n}(z)=\frac{(\hbar^{n/2}-\hbar^{-n/2})}{1-(z_{0}\ldots z_{l-1})^{-n}}$
@@
Note that for $l=1$ we recover the fusion operator for $\text{Hilb}^{n}(\mathbb{C}^{2})$, see section 7 of [48].
@@ 7.1 Limit of vertex function
Let $\mathcal{M}=\mathcal{M}(\mathsf{v},\mathsf{w})$ be a $\hat{A}_{l-1}$ quiver variety. Split the framing as $\mathsf{w}=u_{1}\mathsf{w}^{\prime}+u_{2}\mathsf{w}^{\prime\prime}$ and let the corresponding subtorus of the framing torus be $\mathsf{A}^{\prime}$. We consider the limit of the vertex $V^{\left(\tau\right)}_{\mathcal{M}}$ of $\mathcal{M}$ as $u_{2}\to\infty$. It is proved in Lemma 7.3.5 of [46] that $\lim_{u_{2}\to\infty}V^{\left(\tau\right)}_{\mathcal{M}}$ factorizes, up to a shift, as long as the limit of the descendant $\tau$ exists. Here we make the shift explicit.
@@
We denote $F^{\text{st}}=\mathcal{M}^{\text{st}}(\mathsf{v}^{\prime},\mathsf{w}^{\prime})\times\mathcal{M}^{\text{st}}(\mathsf{v}^{\prime\prime},\mathsf{w}^{\prime\prime})$. Hence if $\tau\in\mathcal{D}(\mathsf{v},\mathsf{w})$, then $\tau|_{F^{\text{st}}}\in\mathcal{D}(\mathsf{v}^{\prime},\mathsf{w}^{\prime})\otimes\mathcal{D}(\mathsf{v}^{\prime\prime},\mathsf{w}^{\prime\prime})\otimes\mathbb{Q}[u_{1}^{\pm 1},u_{2}^{\pm 1}]$.
@@
This is Lemma 7.3.5 of [46] spelled out for cyclic type $A$ quiver varieties. In general, the powers of $q$ arises from the polarization, and the powers of $\hbar^{1/2}$ are the components of the vector $\mathsf{w}^{\prime\prime}+C\mathsf{v}^{\prime\prime}$ and $\mathsf{w}^{\prime}-C\mathsf{v}^{\prime}$, where $C$ is the Cartan matrix of the quiver. ∎
@@
If, in particular, $\mathcal{M}=\mathcal{M}((n,n,\ldots,n),2\delta_{0})$, $\mathcal{M}^{\prime}=\mathcal{M}(n)$, and $\mathcal{M}^{\prime\prime}=\mathcal{M}(0)=\{pt\}$, then
@@ 7.2 Limit of capping operator
Consider $\mathcal{M}:=\mathcal{M}(\mathsf{w})$ and split the framing as $\mathsf{w}=u_{1}\mathsf{w}^{\prime}+u_{2}\mathsf{w}^{\prime\prime}$. Let the corresponding subtorus of the framing torus be $\mathsf{A}^{\prime}$. Then $\mathcal{M}(\mathsf{w})^{\mathsf{A}^{\prime}}=\mathcal{M}(\mathsf{w}^{\prime})\times\mathcal{M}(\mathsf{w}^{\prime\prime})=:\mathcal{M}^{\prime}\times\mathcal{M}^{\prime\prime}$. Let $\text{Stab}:K_{\mathsf{T}}(\mathcal{M}(\mathsf{w})^{\mathsf{A}^{\prime}})\to K_{\mathsf{T}}(\mathcal{M}(\mathsf{w}))$ be the stable envelope [46] with the chamber such that $a=u_{1}/u_{2}$ is an attracting weight, polarization given by
@@
Smirnov’s paper [59] considers only the Jordan quiver. Nevertheless, his techniques apply equally well to all quivers, with the only needed modification being the fixed-component dependent shift in the statement of the theorem.
@@
This follows immediately from the three axioms for stable envelopes, see [48]. ∎
@@
Let $\mathscr{L}_{0}^{(2)}$ be the operator of tensor multiplication by the line bundle $\det\mathcal{V}_{0}$ on $\mathcal{M}(2\delta_{0})$. The component of
@@ 8 Derivation of main formula
We now have all the ingredients needed to prove the main theorem. Recall the quantities $N_{\lambda}$ and $C^{(i)}_{m,n}$ from section 2.
@@
We have
$\sum_{\begin{subarray}{c}\lambda\\ \text{core}_{l}(\lambda)=\emptyset\end{subarray}}N_{\lambda}^{-1}H_{\lambda}\hat{V}^{\left(\tau_{0}\right)}_{\lambda}=\\ \exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}p^{(i)}_{n}}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left(1-\frac{(-\hbar^{1/2}u)^{n}-(\hbar uz_{0}\ldots z_{l-1})^{n}}{(\hbar^{1/2})^{n}-(-z_{0}\ldots z_{l-1})^{n}}\right)\right)$
@@
We first set up the notations for the proof. Let $\mathcal{M}^{(2)}_{\emptyset}=\bigsqcup_{n}\mathcal{M}((n,n,\ldots,n),2\delta_{0})$ and let $\mathsf{A}$ be the (rank two) framing torus, which has coordinates $u_{1}$ and $u_{2}$. The superscript $(2)$ refers to the size of the framing, while the subscript $\emptyset$ refers to the fact that the $K$-theory of $\mathcal{M}^{(2)}_{\emptyset}$ has a basis given by wreath Macdonald polynomials with empty l-core. Let $\tau^{(2)}_{0}$ denote the descendant from Definition 3.4 on $\mathcal{M}^{(2)}_{\emptyset}$ corresponding to the 0th tautological bundle. Let $V^{\left(\tau^{(2)}_{0}\right)}$ denote the corresponding descendant vertex function. Let $\Psi^{(2)}$ denote the capping operator for $\mathcal{M}^{(2)}_{\emptyset}$. Let $\mathscr{L}_{0}^{(2)}$ denote the operator of tensor multiplication in $K$-theory by the 0th tautological line bundle of $\mathcal{M}^{(2)}_{\emptyset}$.
@@
[…]
@@
$\sum_{\begin{subarray}{c}\lambda\\ \text{core}_{l}(\lambda)=\emptyset\end{subarray}}N_{\lambda}^{-1}H_{\lambda}[x]\hat{V}^{\left(\tau_{0}\right)}_{\lambda}|_{z_{0}=z_{0}(-\hbar^{1/2})}\\ =\left(J(z)^{-1}\sum_{\begin{subarray}{c}\lambda,\mu\\ \text{core}_{l}(\lambda)=\emptyset\\ \text{core}_{l}(\mu)=\emptyset\end{subarray}}N_{\lambda}^{-1}N_{\mu}^{-1}H_{\lambda}[x]H_{\mu}[-u_{1}\hbar^{1/2}y]\tau_{0}|_{\lambda}\right)\bigg|_{y=0}$ (19)
From Theorem 6.10, $J(z)^{-1}|_{K_{\mathsf{T}}(\mathcal{M}^{(1)}_{\emptyset})\otimes K_{\mathsf{T}}(\mathcal{M}^{(1)}_{\emptyset})}$, followed by $y=0$, acts by $p^{(i)}_{n}[x]\mapsto p^{(i)}_{n}[x]$ and $p^{(i)}_{n}[y]\mapsto\frac{(\hbar^{-n/2}-\hbar^{n/2})}{1-(z_{0}\ldots z_{l-1})^{-n}}p^{(i)}_{n}[x]$. Combining this with Lemma 2.8 shows that the right side of (19) is equal to
$\left(J(z)^{-1}\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left((1-(-u)^{n})p^{(i)}_{n}[x]+(-u\hbar^{1/2})^{n}p^{(i)}_{n}[y]\right)\right)\right)\Bigg|_{y=0}\\ =\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}p^{(i)}_{n}[x]}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left((1-(-u)^{n})+\frac{(-u)^{n}(1-\hbar^{n})}{1-(z_{0}\ldots z_{l-1})^{-n}}\right)\right)$
@@
$\sum_{\begin{subarray}{c}\lambda\\ \text{core}_{l}(\lambda)=\emptyset\end{subarray}}N_{\lambda}^{-1}H_{\lambda}[x]\hat{V}^{\left(\tau_{0}\right)}_{\lambda}=\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}\frac{C^{(i)}_{0,n}p^{(i)}_{n}[x]}{n(1-t_{1}^{nl})(1-t_{2}^{nl})}\left(1-\frac{(-\hbar^{1/2}u)^{n}-(\hbar uz_{0}\ldots z_{l-1})^{n}}{(\hbar^{1/2})^{n}-(-z_{0}\ldots z_{l-1})^{n}}\right)\right)$